# CommonPower: Supercharging Machine Learning for Smart Grids

### v1

@@ Abstract
The growing complexity of power system management has led to an increased interest in the use of reinforcement learning (RL). However, no tool for comprehensive and realistic benchmarking of RL in smart grids exists. One prerequisite for such a comparison is a safeguarding mechanism since vanilla RL controllers can not guarantee the satisfaction of system constraints. Other central requirements include flexible modeling of benchmarking scenarios, credible baselines, and the possibility to investigate the impact of forecast uncertainties. Our Python tool CommonPower is the first modular framework addressing these needs. CommonPower offers a unified interface for single-agent and multi-agent RL training algorithms and includes a built-in model predictive control approach based on a symbolic representation of the system equations. This makes it possible to combine model predictive controllers with RL controllers in the same system. Leveraging the symbolic system model, CommonPower facilitates the study of safeguarding strategies via the flexible formulation of safety layers. Furthermore equipped with a generic forecasting interface, CommonPower constitutes a versatile tool significantly augmenting the exploration of safe RL controllers in smart grids on several dimensions.
@@ I-A Motivation
Safe RL is an active area of research with a large amount of existing literature [8, 9, 10, 11]. Three different safety levels can be distinguished: Encouraged constraint satisfaction, constraint satisfaction with high probability, and guaranteed constraint satisfaction [10]. Due to the critical nature of power systems, guaranteed constraint satisfaction is the required safety level for the real-world deployment of RL controllers. Some approaches require finding Lyapunov functions to prove stability [12] or control barrier function to prove safety [13] – these are very hard to find for non-experts, especially when the system dynamics is complex.
@@
While CommonPower primarily targets the research community, potential real-world applications are numerous: For the single-agent/centralized set-up, stakeholders could be companies providing building management systems or virtual power plant operators. For competitive or mixed competitive/cooperative scenarios, potential stakeholders are prosumers organized in an energy community or participants of a local energy market.
@@ I-B Contributions
With CommonPower, we provide a platform to accelerate development and enhance comparability within the safe RL community. The specific contributions are as follows.
  • Modular architecture: CommonPower has a highly modular approach in which power system entities, controllers, safety layers, and forecasters are abstracted as objects with clearly defined interfaces.
  • Unified interface for single-agent and multi-agent RL: Contrary to existing tools that provide an interface either for single-agent or multi-agent RL, CommonPower implements a unified interface. This facilitates comparing both approaches.
  • Symbolic system representation: CommonPower comprises two abstract domains. In the object domain, a power system is represented as a tree of objects that can be interacted with. In the symbolic domain this object tree is represented symbolically utilizing the well-established Python optimization framework Pyomo [18, 19]. The symbolic model is central for enabling both the RL safeguarding approach and the built-in optimal control baseline mentioned below.
  • Adaptive RL safeguarding mechanisms: Since CommonPower maintains a symbolic representation of the system under study, a variety of model-based RL safeguarding approaches can be employed by defining safety shields. The default implementation is based on a model predictive control formulation of the system constraints [20], which is used for a safe action projection via optimal programming. This procedure can be applied to arbitrary component models and readily covers single-agent as well as multi-agent setups.
  • Heterogeneous controllers: Due to the modular design of CommonPower, different types of controllers can be combined in the same multi-agent system. To the best of our knowledge, CommonPower is the first tool enabling control schemes with both RL-based and non-RL controllers in the smart grid context.
  • Built-in central/decentral MPC: Evaluating an RL controller against a realistic optimal controller baseline is quite cumbersome in existing frameworks. In contrast to this, CommonPower utilizes its symbolic representation to provide a built-in model predictive controller that can serve as a credible baseline.
  • Generic interfaces for diverse data sources and forecasting models: Existing tools rarely allow studying the influence of inaccurate forecasts of quantities such as loads, electricity prices, or renewable generation on controller performance. CommonPower, however, implements comprehensive interfaces facilitating the use of arbitrary data sources and forecasting models.
A comparison with existing libraries that provide interfaces to RL-based controllers can be found in Tab. I. It shows that CommonPower is the only tool in the smart grid context that provides flexible component modeling, symbolic equations, decentralized control, and flexible prediction models.
@@
The type of setup that can unprecedentedly be modeled in CommonPower is illustrated in Fig. 1. The figure shows decentral control with heterogeneous controllers with safeguarding based on symbolic representations.
Figure 1: Illustration of a decentral control set-up featuring non-RL and RL-based controllers with modular safety layers.
@@ I-C Organization
After discussing related work (Sec. II), we introduce the high-level problem formulation and establish some theoretical RL background (Sec. III). We proceed by describing how power systems (Sec. IV) and control flows (Sec. V) are modeled in CommonPower. Finally, we investigate an exemplary use case (Sec. VI) and conclude (Sec. VII).
@@ II-A Control Applications Utilizing Safe Single-Agent RL
However, the study in [29] argues that such projections might have unfavorable implications on the learning performance of the RL agent, which demands further investigation. A step in this direction is the incorporation of a differentiable projection layer within the policy network of the RL agent [30, 31]. In a similar fashion, the approach in [32] adjusts the output of the policy network in two steps. The first completes the partial network output to satisfy equality constraints, while the second computes a correction to additionally satisfy inequality constraints. In contrast, the work in [33] develops a differentiable function that maps the action of the RL agent to a previously computed robust control invariant set. A different route is taken in [34], where the learned action-value function is represented as a mixed-integer optimal program during deployment. The control action is then computed by solving this optimal program, considering all system constraints.
The controller interface in CommonPower provides structured access to the symbolic model of the controllable subsystem. This makes it possible to integrate arbitrary model-based safeguarding strategies into the framework purely by extension.
@@ II-B Control Applications Utilizing Safe Multi-Agent RL
Local constraints, such as the charge limits of independent energy storage systems, can be addressed using the approaches presented in the previous subsection but are commonly neglected in the multi-agent literature. Some works assume underlying backup controllers such as a battery management system that will override unsafe decisions of individual agents [35, 36]. The study in [17] employs a quadratic program to guarantee local constraint satisfaction. However, it only partially investigates how safeguarding affects the MARL training process.
Global constraints whose satisfaction depends on the decisions of all agents are generally modeled as soft constraints: If these constraints are violated, all agents receive a negative reward that encourages them to reduce the violations. In voltage regulation [37, 38] or frequency control [39, 40] problems, MARL is usually employed to improve scalability over single-agent RL by partitioning the action space. The joint goal of all agents is to reduce the deviation from a desired voltage or frequency – this translates to reducing the violation of a soft constraint. During training, reward shaping encourages the RL agents to avoid such violations and re-establish grid stability as fast as possible.
However, such a global reward induces non-stationarity: All agents adjust their policy simultaneously to obtain a higher reward. Since the reward depends on the decisions of all agents, the individual faces a moving target problem [41]. A common remedy for this problem is to use global observations during training, a paradigm that is referred to as centralized training, decentralized execution (CTDE). This paradigm is the standard approach in voltage and frequency control [37, 38, 39, 40] since no privacy concerns have to be respected in this case. The same applies when MARL is used for home energy management tasks where multiple electric appliances within a building are controlled by multiple agents. These agents must coordinate to achieve global goals such as limiting the dependence on the main grid [42] or keeping the indoor temperature within a desired range [43].
@@ III-A Problem Statement
In decentral control, (LABEL:oc_high_level_coalition_constraints) only contains input coupling constraints within coalitions but not between them since we assume that coalitions do not communicate with each other. This makes it very unlikely that the independent dispatch decisions of all coalitions fulfill the input coupling constraint on the network level. Therefore, in practice, we always model at least one balancing asset in decentral control. Balancing assets are dispatched as a second stage by solving (1) with a single coalition $\mathcal{G}^{i}=\mathcal{A}$. Here, the previously computed input trajectories for all coalitions are concatenated in $u^{\mathcal{P}}$, which is treated as uncontrollable input. The input coupling constraint is then given by $u_{t}^{\mathcal{A}}=h^{\mathcal{A}}(x_{t}^{\mathcal{A}},\begin{bmatrix}\hat{d}^{\mathcal{A}}&u_{t}^{\mathcal{P}}\end{bmatrix}^{\top})$.
@@ III-B Theoretical Background
Even though the presented POMG does not model exactly the same problem as (1), the two are closely related. During training, the RL agents learn the policy $\pi^{i}$ with $u^{i}_{t}~{}=~{}\pi^{i}(o^{i}_{t})$ maximizing the expected cumulative discounted reward $\mathbb{E}_{\pi^{i}}[\sum_{k=0}^{T}\gamma^{k}R^{i}_{t+k+1}]$ [58]. However, vanilla RL cannot consider the constraints (LABEL:oc_high_level_coalition_constraints) except for the input limits which are captured by the action space $\mathcal{U}^{i}$. This is why we require a safety shield like the one described in Sec. V-D. Without any safety considerations, the reward is the negative stage cost $R^{i}=-C^{i}$. In Sec. V-D we describe how to modify the reward function to incorporate information about any interventions of the safety shield. We choose the observation of an agent as the concatenation of its state variables and the forecasts of uncontrollable inputs $o^{i}_{t}=\begin{bmatrix}x^{i}_{t}&\hat{d}^{i}_{[t,T]}\end{bmatrix}^{T}$.
@@ III-C Model-based Safeguarding
Corresponding strategies can be divided into preemptive and post-posed approaches. Preemptive methods adjust the action selection process of the RL algorithm itself, whereas post-posed methods adjust the action selected by the agent before applying it to the environment. While preemptive action masking has a better theoretical justification, it poses two drawbacks in practice. First, the implementation is algorithm-specific and often non-trivial, and second, the safe input set must be computed explicitly, which is usually expensive for continuous action spaces [14].
Since CommonPower is designed to be algorithm-agnostic, we initially focus on post-posed approaches, which can be classified into action replacement and action projection [14]. In action replacement, the first step is to verify whether the action proposed by the RL agent violates any constraints. If that is the case, it is replaced by an action from the safe action space. This safe action can be obtained by either solving a constrained optimization problem with a suitable objective function or by explicitly computing the safe action space, e.g., through reachability analysis and sampling from it. Action projection can be considered a special case of replacement and is achieved by computing the safe action via solving a constrained optimization problem with an objective function that minimizes the distance between the proposed and safe action. Action projection is the default safeguarding approach implemented in CommonPower, with the implementation details provided in Sec. V-D.
@@ IV-A2 Symbolic domain
In the symbolic domain, CommonPower uses the optimization framework Pyomo [18, 19]. The reasoning behind choosing Pyomo over a more generic symbolic library is that the resulting models can readily be used in optimization problems. This is required for optimal control and RL safeguarding. Pyomo uses the concepts of variables Var and parameters Param to differentiate between decision variables and model parameters. To enable a high level of model expressiveness, CommonPower defines a range of model element types that each map to either variables or parameters. Accordingly, these types can be thought of as providing metadata to Pyomo. Additionally, there exist types directly corresponding to constraints and sets in Pyomo.
In CommonPower, an entity can be modeled using the following types of model elements:
  • INPUT (Var): Input variable that is set by a controller.
  • DATA (Param): Exogenous input, read from a data provider.
  • STATE (Var): State variable requiring an initial value.
  • VAR (Var): Generic variable not requiring an initial value.
  • CONSTANT (Param): Fixed Parameter. It is either constant across simulation runs or initialized in each run based on a specific logic.
  • COST (Var): Variable specifically representing a cost.
  • CONSTRAINT (Constraint): Input coupling and dynamics functions are defined with this type.
  • SET (Set): Sets can be useful to specify the values a variable can take.
Note that Pyomo can represent arbitrary nonlinear constraints, which makes it possible to model non-convex dynamics or power flow. Specifically, piece-wise linear functions are often useful in power system modeling, e.g., for the approximation of battery dynamics [61]. CommonPower models such occurrences via mixed integer equations and provides a utility that largely automates the generation of corresponding auxiliary constraints (inspired by [62]). This approach generally leads to mixed-integer quadratically constrained programs, which can be solved more efficiently than general nonlinear programs. For illustration, the model of a simple battery storage system is presented in Tab. II. This model introduces a binary charging indicator variable and corresponding constraints based on the big-M method to represent the piece-wise linear dynamics of batteries.
@@
charging indicator $ec$ VAR $ec\in\{0,1\}$
cost $cost$ COST $cost\in\mathbb{R}$
@@ IV-B Global Symbolic Model
After having modeled a scenario in the object domain, the system is represented by an object tree of entities, with every entity defining its individual symbolic model. Before running a simulation, CommonPower automatically aggregates this distributed symbolic representation into a global symbolic model in the form of (1). To this end, the object tree is traversed and all local symbolic models are added to a hierarchical model in which the original tree structure is maintained via Pyomo blocks. This hierarchical structure can readily be used for optimization within Pyomo and defines the scope of optimal controllers and safety layers in a decentral control scenario.
To make the described modeling approach more tangible, we discuss a highly simplified example. The example system consists of two households and a substation under centralized optimal control. The first household owns a load component, while the second household owns a load and a battery. We consider only active power $p$ and, accordingly, enforce only active power balance on the network level. The battery has a state of charge $soc$ that describes the amount of stored energy. We regard the substation as a trading node with a time-varying unit price $\phi_{t}\geq 0$ and write the set of a node’s children as $\mathcal{W}$. For ease of exposition, we ignore state and input limits. To simulate this system in CommonPower, the user would compose an object tree representing the system structure from instances of predefined or self-modeled entity classes. The resulting representation of the described example system in the object domain is visualized in Fig. 3. Here, each entity instance displays its respective simplified symbolic model.
Figure 3: Example system in the object domain with each entity displaying its symbolic model.
In the example system, the only control input is the active power set point of the battery. The control task is to choose it such that the cost at the substation is minimized under uncertain forecasts for the active power demand of the two loads and the market prices. CommonPower then obtains the global symbolic model for the above example system as
$\begin{split}&\min\sum_{t=0}^{T}\hat{\phi}_{t}\,p^{S1}\\ &s.t.\\ &p_{t}^{H1}=\hat{p}_{t}^{L11}\\ &p_{t}^{H2}=\hat{p}_{t}^{L21}+p_{t}^{B22}\\ &0=p_{t}^{H1}+p_{t}^{H2}+p_{t}^{S1}\\ &soc_{t+1}^{B22}=f^{B22}(soc_{t}^{B22},p_{t}^{B22}).\end{split}$
IV-C Data Providers
Exogenous inputs are fed by data providers, which are each composed of a data source and a forecaster. The data source feeds data to the forecaster that returns a forecast of future values. By defining this dual structure and corresponding interfaces, CommonPower provides maximum flexibility with regard to external data sources and forecasting algorithms.
Forecasters return a prediction for the value of a variable for every time step within the control horizon utilizing some algorithm $\Phi$. We write $\hat{d}_{[t_{0},t_{0}+T]}$ to describe the series of predictions for an exogenous input $d$ for the time from $t_{0}$ to $t_{0}+T$ based on the knowledge at time $t_{0}$. A forecaster has access to past and present values of a set of features $z\in\mathcal{Z}$ and optionally to external forecasts of a set of features $\hat{z}\in\hat{\mathcal{Z}}^{\prime}$. Features can be measurements and additional information, such as the day of the week. A forecaster then implements
$\hat{d}_{[t_{0},t_{0}+T]}=\Phi(z_{[t_{0}-L,t_{0}-1]},\hat{z}^{\prime}_{[t_{0},t_{0}+T]}),$ (2)
where $L$ is the number of time steps to look back into the past. Note that we do not generally require that $d\in\mathcal{Z}$.

### v3

# CommonPower: A Framework for Safe Data-Driven Smart Grid Control

@@ Abstract
The growing complexity of power system management has led to an increased interest in reinforcement learning (RL). However, vanilla RL controllers cannot themselves ensure satisfaction of system constraints. Therefore, combining them with formally correct safeguarding mechanisms is an important aspect when studying RL for power system management. Integrating safeguarding into complex use cases requires tool support. To address this need, we introduce the Python tool CommonPower. CommonPower’s unique contribution lies in its symbolic modeling approach, which enables flexible, model-based safeguarding of RL controllers. Moreover, CommonPower offers a unified interface for single-agent RL, multi-agent RL, and optimal control, with seamless integration of different forecasting methods. This allows users to validate the effectiveness of safe RL controllers across a large variety of case studies and investigate the influence of specific aspects on overall performance. We demonstrate CommonPower’s versatility through a numerical case study that compares RL agents featuring different safeguards with a model predictive controller in the context of building energy management.
@@ I-A Motivation
Safe RL is an active area of research with a large amount of existing literature [8, 9, 10, 11]. Three different safety levels can be distinguished: Encouraged constraint satisfaction, constraint satisfaction with high probability, and guaranteed constraint satisfaction [10]. Due to the critical nature of power systems, guaranteed constraint satisfaction is the required safety level for the real-world deployment of RL controllers. Some approaches require finding Lyapunov functions to prove stability [12] or control barrier functions to prove safety [13] – these are very hard to find for non-experts, especially when the system dynamics are complex.
@@ I-B Contributions
The central aim of CommonPower is to pave the way toward comprehensive analysis of safe RL-based controllers in the smart grid context. This work introduces the required theoretical abstractions, describes the corresponding implementation in CommonPower, and reports a case study on the influence of different safeguarding approaches. Specifically, our main contributions are:
  • Formally correct model-based safeguarding: The presence of inaccurate forecasts must be explicitly considered in the safeguarding mechanisms. We provide a general formalization for model-based safeguarding robustified with approaches from reachability analysis.
  • Control on arbitrary levels of decentralization: A coalition-based formulation enables unified treatment of single-agent and multi-agent RL and the combination of RL-based and non-RL controllers in the same system.
  • Modular Python library CommonPower: CommonPower facilitates component modeling via symbolic representations. These are used to automatically extract system constraints for safeguarding and a built-in model predictive controller. Furthermore, the library implements a unified interface for single-agent and multi-agent RL algorithms via a gymnasium interface [18]. Due to the highly modular structure, custom controllers, safeguarding mechanisms, and forecasters can be integrated purely by extension.
  • Numerical benchmark of different safeguarding approaches: We utilize CommonPower to show how different safeguarding types affect the performance of an RL agent tasked with the optimal scheduling of a battery.
Existing libraries that provide interfaces to RL-based controllers do not provide the modularity that would facilitate straightforward investigations along the above-mentioned dimensions. The comparison in Tab. I shows that CommonPower is the only tool in the smart grid context that provides flexible component modeling, symbolic equations, decentralized control, and flexible prediction models.
@@
The type of setup that can unprecedentedly be modeled in CommonPower is illustrated in Fig. 1. The figure shows decentral control with heterogeneous controllers and safeguarding based on symbolic representations.
Figure 1: Illustration of a decentral control set-up featuring non-RL and RL-based controllers with modular safeguards.
@@ I-C Organization
After discussing related work (Sec. II), we introduce the high-level problem formulation and establish some theoretical RL background (Sec. III). We describe how power systems and forecasters are modeled in CommonPower (Sec. IV) and present our unified approaches to control and safeguarding (Sec. V). Finally, we investigate an exemplary case study (Sec. VI) and conclude (Sec. VII).
@@ II-A Control Applications Utilizing Safe Single-Agent RL
However, the study in [27] argues that such projections might have unfavorable implications on the learning performance of the RL agent, which demands further investigation. A step in this direction is the incorporation of a differentiable projection layer within the policy network of the RL agent [28, 29]. In a similar fashion, the approach in [30] adjusts the output of the policy network in two steps. The first completes the partial network output to satisfy equality constraints, while the second computes a correction to additionally satisfy inequality constraints. In contrast, the work in [31] develops a differentiable function that maps the action of the RL agent to a previously computed robust control invariant set.
@@ II-B Control Applications Utilizing Safe Multi-Agent RL
Local constraints, such as the charge limits of independent energy storage systems, can be addressed using the approaches presented in the previous subsection but are commonly neglected in the multi-agent literature. Some works assume underlying backup controllers such as a battery management system that will override unsafe decisions of individual agents [32, 33]. The study in [17] employs a quadratic program to guarantee local constraint satisfaction.
Global constraints whose satisfaction depends on the decisions of all agents are generally modeled as soft constraints: If these constraints are violated, all agents receive a negative reward encouraging them to reduce violations. In voltage regulation [34, 35] or frequency control [36, 37], a deviation from the desired voltage or frequency will be penalized. When using MARL in the context of home energy management tasks where multiple electrical appliances within a building are controlled by multiple agents, penalties are used to reduce dependence on the main grid [38] or maintain comfort [39].
However, such a global reward induces non-stationarity: All agents adjust their policy simultaneously to obtain a higher reward. Since the reward depends on the decisions of all agents, the individual faces a moving target problem [40]. A common remedy for this problem is to use global observations during training, a paradigm that is referred to as centralized training, decentralized execution (CTDE).
@@ III-A Problem Statement
In decentral control, (LABEL:oc_high_level_coalition_constraints) only contains input coupling constraints within coalitions but not between them since we assume that coalitions do not communicate with each other. This means that the independent dispatch decisions of all coalitions might not fulfill all network-level constraints. Therefore, in practice, we always model at least one balancing asset in decentral control. Balancing assets are dispatched as a second stage by solving (1) with a single coalition $\mathcal{G}^{i}=\mathcal{A}$. Here, the previously computed input trajectories for all coalitions are concatenated in $u^{\mathcal{P}}$, which is treated as uncontrollable input. The input coupling constraint is then given by $u_{t}^{\mathcal{A}}=h^{\mathcal{A}}(x_{t}^{\mathcal{A}},\begin{bmatrix}\hat{d}^{\mathcal{A}}&u_{t}^{\mathcal{P}}\end{bmatrix}^{\top})$.
@@ III-B Theoretical Background
Even though the presented POMG does not model exactly the same problem as (1), the two are closely related. During training, the RL agents learn the policy $\pi^{i}$ with $u^{i}_{t}~{}=~{}\pi^{i}(o^{i}_{t})$ maximizing the expected cumulative discounted reward $\mathbb{E}_{\pi^{i}}[\sum_{k=0}^{T}\gamma^{k}R^{i}_{t+k+1}]$ [55]. However, vanilla RL cannot consider the constraints (LABEL:oc_high_level_coalition_constraints) except for the input limits which are captured by the action space $\mathcal{U}^{i}$. This is why we require a safeguard like the one described in Sec. V-B. Without any safety considerations, the reward is the negative stage cost $R^{i}=-C^{i}$. In Sec. V-B we describe how to modify the reward function to incorporate information about any interventions of the safeguard. We choose the observation of an agent as the concatenation of its state variables and the forecasts of uncontrollable inputs $o^{i}_{t}=\begin{bmatrix}x^{i}_{t}&\hat{d}^{i}_{[t,T]}\end{bmatrix}^{T}$.
@@ III-C Model-based Safeguarding
Corresponding strategies can be divided into preemptive and post-posed approaches. Preemptive methods adjust the action selection process of the RL algorithm itself, whereas post-posed methods adjust the action selected by the agent before applying it to the environment. Since our framework is algorithm-agnostic, we only provide an abstraction and corresponding implementation for post-posed approaches, as detailed in Sec. V-B.
@@ IV-A2 Symbolic domain
In the symbolic domain, CommonPower uses the optimization framework Pyomo [59, 60]. The reasoning behind choosing Pyomo over a more generic symbolic library is that the resulting models can readily be used in optimization problems. This is required for optimal control and RL safeguarding. Pyomo uses the concepts of variables Var and parameters Param to differentiate between decision variables and model parameters. To enable a high level of model expressiveness, CommonPower defines a range of model element types that each map to either variables or parameters. Accordingly, these types can be thought of as providing metadata to Pyomo. Additionally, there exist types directly corresponding to constraints and sets in Pyomo. Note that Pyomo can represent arbitrary nonlinear constraints, which makes it possible to model non-convex dynamics or power flow. Specifically, piece-wise linear functions are often useful in power system modeling, e.g., for the approximation of battery dynamics [61]. CommonPower models such occurrences via mixed integer equations and provides a utility that largely automates the generation of corresponding auxiliary constraints (inspired by [62]). This approach generally leads to mixed-integer quadratically constrained programs, which can be solved more efficiently than general nonlinear programs. For illustration, the model of a simple battery storage system is presented in Tab. II. This model introduces a binary charging indicator variable and corresponding constraints based on the big-M method to represent the piece-wise linear dynamics of batteries.
@@
charging indicator $ec$ state $ec\in\{0,1\}$
cost $cost$ state $cost\in\mathbb{R}$
@@ IV-B Global Symbolic Model
After having modeled a scenario in the object domain, the system is represented by an object tree of entities, with every entity defining its individual symbolic model. Before running a simulation, CommonPower automatically aggregates this distributed symbolic representation into a global symbolic model in the form of (1). To this end, the object tree is traversed and all local symbolic models are added to a hierarchical model in which the original tree structure is maintained via Pyomo blocks. This hierarchical structure can readily be used for optimization within Pyomo and defines the scope of optimal controllers and safeguards in a decentral control scenario.
IV-C Forecasters
CommonPower provides maximum flexibility by defining high-level interfaces for data sources and forecasting algorithms. Forecasters return a prediction for the value of a variable for every time step within the control horizon while having access to past and present values of a set of features. Features can be measurements and additional information, such as the day of the week. In economic dispatch, the quality of forecasts significantly influences the final system cost. To enable an explicit treatment of the associated uncertainty in the safeguarding mechanism, forecasters can return forecast bounds. These bounds signify upper limits $\bar{d}_{t}$ and lower limits $\underline{d}_{t}$, which we treat as guarantees obtained by conformance checking [63]. By default, we assume guaranteed least-conservative bounds, i.e., bounds based on the absolute difference between forecast and the true value, formally $\hat{d}\in[d_{t}+|d_{t}-\hat{d_{t}}|,d_{t}-|d_{t}-\hat{d_{t}}|]$.
CommonPower implements a framework for training, evaluating, and tuning machine learning models to facilitate the study of data-driven forecasting approaches. To obtain a trained forecasting model on a given data source, the user only needs to provide a custom model that implements PyTorch’s nn.Module interface [64] and some parameters for the training pipeline.